Suppose that the moment generating function of X is M(t) 1-2t . Find E[X]rounded to nearest .xx.
Answer #1

We know the following relation between the expected value of a random variable and its moment generating function:

\begin{align*} \boldsymbol{E[X]} &= \left.\frac{\mathrm{d} M(t)}{\mathrm{d} t} \right |_{t=0} \\ &= \left.\frac{\mathrm{d} }{\mathrm{d} t}\left(\frac{1}{1-2t} \right ) \right |_{t=0} \\ &= \left.\frac{\mathrm{d} (1-2t)^{-1}}{\mathrm{d} t} \right |_{t=0} \\ &= \left. \left[(-1)*(1-2t)^{-1-1}* \frac{\mathrm{d} (1-2t)}{\mathrm{d} t}\right ]\right |_{t=0} \\ &\left[\text{Using Chain Rule} \right ] \\ &= \left. \left[-(1-2t)^{-2}* (-2) \right ]\right |_{t=0} \\ &= \left. \frac{2}{(1-2t)^2}\right |_{t=0} \\ &= \frac{2}{(1-2*0)^2} \\ &= \frac{2}{(1-0)^2} \\ &= \frac{2}{1^2} \\&= \bf 2.00 \ \ \ \ \ \ \ \ \ \ [ANSWER] \end{align*}

For any queries, feel free to comment and ask.

If the solution was helpful to you, don't forget to upvote it by clicking on the 'thumbs up' button.